# Superconducting pairing correlations on a trapped-ion quantum computer

### v1

@@ V Conclusion
$U\approx\exp\left(\frac{U\tau}{2}\sum_{i}n_{i\uparrow}n_{i\downarrow}\right)$ (S85)
$\times\exp\left(-t\tau\sum_{\langle ij\rangle,\sigma}\alpha_{\langle ij\rangle}c_{i\sigma}^{\dagger}c_{j\sigma}\right)\exp\left(\frac{U\tau}{2}\sum_{i}n_{i\uparrow}n_{i\downarrow}\right)\,$,
@@
We presented in the main text a benchmark of the hardware for the simulation of fermionic systems. The benchmark relied on the case $U=0$ where there is no interaction between the fermions, for which the imbalance can be computed in polynomial time. Such benchmark protocol was defined in Ref. [58]. There, was introduced the notion of distinguishability cost to evaluate the quality of the output of a quantum hardware. This distinguishability cost is defined as the minimal amount of resources that an ideal perfect quantum computer has to spent to certify that the output of the tested hardware is incorrect. This amount of resources can be measured in shots or in number of gates to implement for example. In case of just one output $m$ with standard deviation $\sigma$ obtained from the hardware, and ideal value $m_{0}$ with standard deviation per shot $\sigma_{0}$ on the perfect hardware, this distinguishability cost is defined as
@@
where $N_{\rm gates}$ is the number of two-qubit gates in the circuit tested, and $\lceil\cdot\rceil$ the rounding to the nearest larger integer. This quantity $\delta$ corresponds exactly to the total number of two-qubit gates to implement (across different shots) on a perfect quantum computer to be confident by three standard deviations that the outcome of the tested hardware, whose precision is bounded by $\sigma$, is incorrect. The exact values $m_{0}$ and the variance per shot $\sigma_{0}$ can be computed exactly for this free-fermion benchmark. In case of multiple time points, the definition of the distinguishability cost is more involved and requires a $\chi^{2}$ test, see [58]. The interpretation of this score is that before running $\delta$ two-qubit gates, the outcome of the tested hardware is statistically compatible with that of a perfect hardware. The score automatically balances shot noise and hardware noise and puts them on same footing.
@@
with $\sigma=1$. This filters the sum to only terms located at distance $d$ in Manhattan distance. We present the results in Fig S20c. We observe that as expected, the staggered $\eta$-pairing is strongly short-ranged right after the pulse, whereas after relaxation it vanishes at $d=0$, peaks at $d=2.3$ and has non-zero tails at larger distance. This increase of long-distance staggered $\eta$-pairing correlations after relaxation was also observed in a one-dimensional setting [30].
@@
The key question is then: how difficult, in terms of bond dimension is it to achieve order-one fidelities using the circuit-DMRG method, for the larger $6\times 6$ system? To answer this question, we show simulation fidelity as a function of bond dimension in Fig. S22b. Even the shallowest (2 steps) circuit achieves a fidelity roughly 3000 times lower that of the hardest Ising quench considered in [60] at bond dimension $\chi=2048$.
@@
The extremely small fidelity achieved for these circuits is not very surprising: A significant fraction of two-qubit gates are Clifford gates, most of which are used to couple the system qubits to the ancilla qubits during the fermionic hopping gadget. While these are necessary to enforce the fermion-to-qubit mapping, classical methods can achieve the correct fermionic statistics in other ways. For example, one can associate the physical degree of freedom of each MPS tensor to a fermionic mode and enforce the fermionic parity by introducing a $Z_{2}$ symmetry [62]. To lower bound the cost of such an approach, we have also considered circuits in which all ancilla qubits (and gates acting on them) are removed, as well as all other gates that are responsible for the fermion-to-qubit encoding. The resulting simulation fidelities, shown in Fig. S22c, are much larger than those where the ancilla degrees of freedom are kept track of explicitly. In particular, the shallowest (2 steps) circuit now becomes comparable in difficulty to the $s=20$ Ising quench discussed above. As a result, we expect that a fermionic circuit-DMRG method utilizing additional symmetries, for example $SU(2)\times U(1)$ or $U(1)\times U(1)$ over spin and number of particles, may be able to capture the nearest-neighbour correlations accurately [63]. The situation on the deeper circuits (3 steps and 4 steps), where after the light pulse additional Hubbard-Trotter steps are applied to relax the state, remain out of reach.
@@
In addition to the circuit-DMRG results obtained from the FermioniQ Ava emulator, we have also run some simulations of the circuit written with the Jordan-Wigner (JW) encoding, alleviating the difficulty of simulating the compact encoding, at the price of adding long range JW strings. The JW string order is chosen to match the ordering of MPS tensors, making this approach similar to the fermionic MPS approach without having to utilize an extra $Z_{2}$ symmetry. These MPS simulations were performed using a slightly simpler approach where quantum gates are contracted into the MPS one at a time. The algorithm supports native application of two-qubit gates between non-adjacent qubits, as well as multi-qubit gates of the form $\exp(i\theta P)$ for arbitrary Pauli strings. These are applied by first obtaining their exact representation as an MPO of bond dimension 2, which is then contracted into the MPS following the zip-up algorithm from [64], with one caveat: the zip-up algorithm is carried out using QR decompositions, rather than SVD. After the gate has been contracted into the MPS, the latter is put in canonical form with the center at the leftmost bond affected by the gate, followed by SVD truncation down to the target bond dimension; this is repeated on all affected bonds from left to right. As discussed in [65], the sum of the squared singular values we keep determines the fidelity of each single truncation, and the product of all of these is a standard estimate of the fidelity of the final state. The algorithm runs on GPUs, implemented using NVIDIA cuTensorNet and is publicly available at [66] under the name MPSxGate. The results can be seen (blue crosses) in Fig. S22c. Interestingly, the fidelity match closely the circuit simulations where all ancilla have been removed. This indicates the naive estimate by removing gates related to ancilla gives a reasonable estimate of the complexity of simulating with fermionic MPS.
@@
Figure S22: Circuit Simulation with Fermioniq’s Ava circuit-DMRG emulator on the light-matter interaction circuits in Fig. 1g. (a) Verification on the exactly solvable $4\times 4$ case, after application of the 2-Trotter step light pulse, before any further relaxation. The exact nearest-neighbour $\Delta_{\eta}$ correlation is obtained by translating the circuit to a 32-qubit Jordan-Wigner circuit and computing the exact statevector result. Compatible with the findings in [60], (extrapolated) results from circuit-DMRG are reliable when a simulation fidelity $F$ of order 1 can be achieved, and are essentially random for $F\ll 50\%$. (b) Simulation fidelities on the $6\times 6$ circuits on 90 qubits, right after the light pulse (2 steps) and after up to two Hubbard-Trotter relaxation steps (3 and 4 steps), compared with the “intermediate temperature” ($\Delta\theta=2\pi/9$) quench circuit with $s=20$ Ising-Trotter steps in [60]. (c) Circuit fidelities on 72-qubit circuits that are obtained by deleting all gates on the ancilla qubits, as well as any maximum angle CZZ gates, which are responsible for implementing the fermion-to-qubit encoding. The result is a lower bound on the cost of natively fermionic circuit-DMRG.
@@
We now study the simulability of the circuits with Sparse Pauli Decomposition (SPD). This is a promising simulation technique that has been attracting attention recently [67, 68, 69, 70, 71] and which we have found most practical for the development and testing of quantum circuits in this work. Indeed, it displays a fast and precise convergence for observables with few enough gates in their light-cone for an arbitrary number of qubits (and in general, for circuits with finite magic [72, 73]). This simulation technique works in the Heisenberg picture, and consists in expanding the observable to be measured into strings of Pauli matrices
@@
A generic operator to be decomposed as in (S109) would require $4^{N}$ terms, which becomes intractable for $N\gtrapprox 20$. Starting from a local observable that involves only a few Pauli matrices, we can keep perform the simulation by storing only the non-zero terms in the decomposition (S111). However, the number of gates to keep track of grows exponentially (generically, is multiplied by $2$) after every gate application. The exact bookkeeping of all the non-zero terms ends up thus being unfeasible. As a result, different approximations can be made to get a certain precision on the result. The approximation that we will consider consist in systematically truncating, after every application of gates, all the terms $c_{n}$ that are below some threshold $\epsilon$ in absolute value. Namely, one retains in the sum (S111) only the terms $c_{n}$ such that $|c_{n}|\geq\epsilon$. The complexity of the simulation is related to the non-stabilizerness of the operator, which can be quantified by the operator stabilizer entropy [72]. This simulation technique is particularly appealing for finite-magic circuits and for code testing, as it is not limited by the number of qubits and the circuit geometry. This truncation scheme has been used in the largest-scale SPD numerical simulation to date [68].
@@
In this resource estimation study, we will first focus on the memory requirement, namely on the number of Pauli strings required in the SPD to obtain a certain precision on the observable. We will then discuss an estimation of the actual runtime based on comparison with previous works. We note that Majorana fermion decomposition (see Ref. [74]) instead of Pauli string decompositions are not expected to give an advantage in terms of memory, because within the fermionic encoding there should be a one-to-one correspondence between Pauli strings and Majorana strings.
@@
This is a rather standard metric for estimating the difficulty of the simulation, that has been employed before [70]. However, in our case, the circuit has a particular structure. Before the “injection”, the gates at the beginning of the circuit only act on the first $L=36$ qubits, for $90$ qubits in total. This means that in the Heisenberg picture, after the injection is passed, one can keep in the SPD only the strings that have either $I$ or $Z$ Pauli matrices on all the remaining $54$ qubits. This significantly reduces the number of strings to keep track of, but this also modifies the norm, because this projection is non-unitary. Since we would like to know the norm at the end of the circuit, we do not perform this projection. Therefore, for a fairer resource estimate, we will consider the maximal number of strings only before (in the Heisenberg picture) the injection, namely
@@
Figure S25: Relative norm squared as a function of maximal number of strings before injection $\mathcal{S}_{\rm metric}$. Data shown is for the circuits run on hardware on a $6\times 6$ lattice, performing 22 Trotter steps after state preparation (corresponding to the light pulse), 33 Trotter steps (one extra step on top of the light pulse), and 44 Trotter steps (two extra steps). The dashed lines indicate quadratic extrapolations using the last 88 points, and dotted lines linear extrapolations using the last 33 points. The gray shaded area indicates the region that is beyond the largest SPD simulation that has been performed before the present work on a supercomputer [68]. The blue shaded area indicates the region where $90\%$ of the norm is obtained.
@@
How do these numbers translate to actual runtimes? This question is generally difficult to answer due to differences in classical compute hardware and implementation, although in this case we can make some estimates based on the recent parallelised implementation of SPD on the Fugaku supercomputer [68]. In Fig. 3 of that work, wall time per gate is shown for different numbers of Pauli strings $\mathcal{S}$ and parallel processes $M$. Using plot digitisation software [75], we have extracted the scaling of wall time per gate $w$ as a function of $\mathcal{S}$ for a fixed number of parallel processes $M=2^{7}$. While we found the slope $w\sim\mathcal{S}^{1.05}$ to fit the data best, let us make the optimistic assumption that $w\propto\mathcal{S}$ scales linear, that memory is unlimited, and that furthermore the problem can be parallelised perfectly up to $M=2^{22}$ parallel processes. Given that in that work up to 12,288 nodes were used to carry out $2^{18}$ parallel processes, this larger $M$ upper bounds the number of parallel processes that could be run on all 158,976 of Fugaku’s nodes. Under these assumptions, we find a best fit
@@
Figure S26: Estimating the run time for Sparse Pauli Decomposition on Fugaku. (a): Wall time per gate extracted from Fig. 3 in Broers ’25 which refers to Ref. [68] as a function of the number of Pauli strings $\mathcal{S}$. Assuming a perfectly linear scaling of wall time per gate in $\mathcal{S}$ and perfect parallelisation, i.e. inverse linear scaling in the number of parallel processes $M$, the best fit we obtain is $\mathcal{S}/M\times 100$ nanoseconds per gate. (b): Wall time per gate using $M=2^{22}$ parallel processes assuming the scaling extracted in (a) for simulating the 3-step circuit in three different scenarios: A quadratic and linear scenario, as well as the geometric mean, denoted “mid”. We consider the mid scenario to be the most likely of the three and $M^{22}$ to upper bound the maximum number of parallel processes achievable on Fugaku based on Appendix F of [68].
@@
Projected-entangled-pair state (PEPS) methods have become a cornerstone tensor-network framework for simulating quantum many-body systems [62]. In contrast to MPS, PEPS employs a lattice-based architecture that more naturally and efficiently represents entanglement in higher-dimensional ($D>1$) systems. This enhanced expressiveness, however, comes at the cost of greater computational complexity and the absence of a canonical form for optimal compression. We employ a PEPS specifically adapted to the Octagon lattice geometry with periodic boundary conditions (PBC) and integrate a Belief Propagation (BP) algorithm to efficiently evolve the state [76].
@@
For comparison, we also report the fidelities [see Eq. (S108)] using MPS simulations, as well as those obtained using matrix-product operator (MPO) Heisenberg evolution. Notably, MPO evolution proceeds backward in time, and its accuracy depends on both the spatial support of the initial (unevolved) operator and the light-cone structure of the circuit. The fidelity of MPO simulations drops rapidly as the light cone of the operator expands, indicating that MPO methods alone are clearly incapable of simulating the full circuit. Similar to both the MPS and SPD approach, linear extrapolation of $\langle O(F)\rangle$ using MPOs has also been considered to extend the limit of finite bond dimension. Empirically, fidelity-based extrapolation of expectation values has been reported to yield results consistent with exact values in the IBM quantum utility experiment [78]. However, the expectation value as a function of fidelity is not monotonic. Even as a heuristic method, the fidelity has to be larger than a certain threshold $\mathcal{F}_{\textrm{th}}$ to obtain meaningful signal, either by extrapolation or claiming convergence of the method. Since the MPO and SPD are similar techniques—both track the evolution of observables in the Heisenberg picture and differ only in their truncation schemes—we expect this threshold to be close to $1$, which makes the stand-alone MPO approach practically infeasible and motivates the use of hybrid techniques.
@@
One possible direction to overcome these limitations is a hybrid approach that partially evolves the state in the Schrödinger picture and the operator in the Heisenberg picture, meeting in the middle — a strategy referred to as the mixed representation. For the $4\times 4$ system, one can imagine evolving MPS and MPO and stopping near where their fidelity curves cross to achieve better fidelity. However, given the near-vanishing MPS fidelity for the $6\times 6$ system, one would instead need to evolve PEPS and MPO (or potentially PEPO) and meet at their crossing point to obtain reasonable fidelities. However, measuring observables (in the form of an MPO or PEPO) for PEPS with $\chi>8$ likely requires either a (potentially sign-problematic) sampling technique or a loop-series BP expansion [79]. It is unclear to us whether the observable can be accurately estimated in such a mixed representation with feasible computing resources, but it would be an interesting direction for further study. Finally, we note that all of these methods could also be implemented with an intrinsically fermionic PEPS and/or PEPO representation that avoids additional overhead of simulating the fermion-to-qubit encoding, generally resulting in lower bond dimension for the relevant objects. Reliably establishing the difficulty of each of these potential classical techniques for simulating these circuits requires significant further study and is left for future work.
@@
The second reason for choosing the mixed boundary conditions is that, ultimately, we are interested in the thermodynamic limit $N\to\infty$ where the boundary conditions have no effects anymore. In that respect, periodic, antiperiodic or mixed boundary conditions are only ways of dealing with the finite system, and are physically equally relevant. One can thus choose the boundary condition that will converge the fastest to the thermodynamic limit. It is known that averaging over “twisted” boundary conditions where a phase $e^{i\varphi}$ is given to fermions that wind around the lattice, with $\varphi$ the value to average over, make expectation values of observables converge faster to the thermodynamic limit than periodic boundary conditions [80]. In that point of view, the mixed boundary conditions can be seen as a rough approximation of this phase averaging, and is thus expected to give results that are closer to the thermodynamic limit than periodic boundary conditions.
@@
In Fig. S28a, we show how different boundary conditions approach the thermodynamic limit. While open boundary conditions (OBC), periodic boundary conditions (PBC) and antiperiodic boundary conditions (in which every hopping term across the boundary comes with an extra minus sign, APBC) systematically underestimate the ground state energy, boundary conditions with PBC in one direction and APBC in the other underestimate it. Mixed periodic boundary conditions (mPBC), which we define by averaging over the 4 settings (PBC-PBC, APBC-APBC, APBC-PBC and PBC-APBC), thus yields the smallest finite-size effects. The resulting relative error with respect to the thermodynamic limit is shown in Fig. S28b. For the $L=6$ lattice considered in this work, mixed periodic boundary conditions already show more than 8 times smaller finite-size effects. To match the finite size effects of a $6\times 6$ mixed periodic lattice, an open boundary conditions setup would require more than $40^{2}=1600$ sites. If such an OBC system was realised using the same fermion-to-qubit mapping as described here, this would increase the qubit count from 90 to $1600\times 2\times 5/4=4000$ qubits. This discrepancy increases with increasing $L$ due to the different slopes in Fig. S28b: While in principle there exist exact expressions for the finite size relative error $r_{L}=|e_{L}-e_{\infty}|/e_{\infty}$, we can obtain a simple fit using the data points from $L=32$ to $L=40$, to obtain

### v2

@@ V Conclusion
$U\approx\exp\left(\frac{iU\tau}{2}\sum_{i}n_{i\uparrow}n_{i\downarrow}\right)$ (S87)
$\times\exp\left(-it\tau\sum_{\langle ij\rangle,\sigma}\alpha_{\langle ij\rangle}c_{i\sigma}^{\dagger}c_{j\sigma}\right)\exp\left(\frac{iU\tau}{2}\sum_{i}n_{i\uparrow}n_{i\downarrow}\right)\,$,
@@
We presented in the main text a benchmark of the hardware for the simulation of fermionic systems. The benchmark relied on the case $U=0$ where there is no interaction between the fermions, for which the imbalance can be computed in polynomial time. Such benchmark protocol was defined in Ref. [granet2025appqsim]. There, was introduced the notion of distinguishability cost to evaluate the quality of the output of a quantum hardware. This distinguishability cost is defined as the minimal amount of resources that an ideal perfect quantum computer has to spent to certify that the output of the tested hardware is incorrect. This amount of resources can be measured in shots or in number of gates to implement for example. In case of just one output $m$ with standard deviation $\sigma$ obtained from the hardware, and ideal value $m_{0}$ with standard deviation per shot $\sigma_{0}$ on the perfect hardware, this distinguishability cost is defined as
@@
where $N_{\rm gates}$ is the number of two-qubit gates in the circuit tested, and $\lceil\cdot\rceil$ the rounding to the nearest larger integer. This quantity $\delta$ corresponds exactly to the total number of two-qubit gates to implement (across different shots) on a perfect quantum computer to be confident by three standard deviations that the outcome of the tested hardware, whose precision is bounded by $\sigma$, is incorrect. The exact values $m_{0}$ and the variance per shot $\sigma_{0}$ can be computed exactly for this free-fermion benchmark. In case of multiple time points, the definition of the distinguishability cost is more involved and requires a $\chi^{2}$ test, see [granet2025appqsim]. The interpretation of this score is that before running $\delta$ two-qubit gates, the outcome of the tested hardware is statistically compatible with that of a perfect hardware. The score automatically balances shot noise and hardware noise and puts them on same footing.
@@
with $\sigma=1$. This filters the sum to only terms located at distance $d$ in Manhattan distance. We present the results in Fig S20c. We observe that as expected, the staggered $\eta$-pairing is strongly short-ranged right after the pulse, whereas after relaxation it vanishes at $d=0$, peaks at $d=2.3$ and has non-zero tails at larger distance. This increase of long-distance staggered $\eta$-pairing correlations after relaxation was also observed in a one-dimensional setting [kaneko_photoinduced_2019].
@@
The key question is then: how difficult, in terms of bond dimension is it to achieve order-one fidelities using the circuit-DMRG method, for the larger $6\times 6$ system? To answer this question, we show simulation fidelity as a function of bond dimension in Fig. S22b. Even the shallowest (2 steps) circuit achieves a fidelity roughly 3000 times lower that of the hardest Ising quench considered in [haghshenas_digital_2025] at bond dimension $\chi=2048$.
@@
The extremely small fidelity achieved for these circuits is not very surprising: A significant fraction of two-qubit gates are Clifford gates, most of which are used to couple the system qubits to the ancilla qubits during the fermionic hopping gadget. While these are necessary to enforce the fermion-to-qubit mapping, classical methods can achieve the correct fermionic statistics in other ways. For example, one can associate the physical degree of freedom of each MPS tensor to a fermionic mode and enforce the fermionic parity by introducing a $Z_{2}$ symmetry [cirac2021matrix]. To lower bound the cost of such an approach, we have also considered circuits in which all ancilla qubits (and gates acting on them) are removed, as well as all other gates that are responsible for the fermion-to-qubit encoding. The resulting simulation fidelities, shown in Fig. S22c, are much larger than those where the ancilla degrees of freedom are kept track of explicitly. In particular, the shallowest (2 steps) circuit now becomes comparable in difficulty to the $s=20$ Ising quench discussed above. As a result, we expect that a fermionic circuit-DMRG method utilizing additional symmetries, for example $SU(2)\times U(1)$ or $U(1)\times U(1)$ over spin and number of particles, may be able to capture the nearest-neighbour correlations accurately [liu2025accurate]. The situation on the deeper circuits (3 steps and 4 steps), where after the light pulse additional Hubbard-Trotter steps are applied to relax the state, remain out of reach.
@@
In addition to the circuit-DMRG results obtained from the FermioniQ Ava emulator, we have also run some simulations of the circuit written with the Jordan-Wigner (JW) encoding, alleviating the difficulty of simulating the compact encoding, at the price of adding long range JW strings. The JW string order is chosen to match the ordering of MPS tensors, making this approach similar to the fermionic MPS approach without having to utilize an extra $Z_{2}$ symmetry. These MPS simulations were performed using a slightly simpler approach where quantum gates are contracted into the MPS one at a time. The algorithm supports native application of two-qubit gates between non-adjacent qubits, as well as multi-qubit gates of the form $\exp(i\theta P)$ for arbitrary Pauli strings. These are applied by first obtaining their exact representation as an MPO of bond dimension 2, which is then contracted into the MPS following the zip-up algorithm from [Stoudenmire_2010], with one caveat: the zip-up algorithm is carried out using QR decompositions, rather than SVD. After the gate has been contracted into the MPS, the latter is put in canonical form with the center at the leftmost bond affected by the gate, followed by SVD truncation down to the target bond dimension; this is repeated on all affected bonds from left to right. As discussed in [Zhou2020], the sum of the squared singular values we keep determines the fidelity of each single truncation, and the product of all of these is a standard estimate of the fidelity of the final state. The algorithm runs on GPUs, implemented using NVIDIA cuTensorNet and is publicly available at [PabloAndresCQ] under the name MPSxGate. The results can be seen (blue crosses) in Fig. S22c. Interestingly, the fidelity match closely the circuit simulations where all ancilla have been removed. This indicates the naive estimate by removing gates related to ancilla gives a reasonable estimate of the complexity of simulating with fermionic MPS.
@@
Figure S22: Circuit Simulation with Fermioniq’s Ava circuit-DMRG emulator on the light-matter interaction circuits in Fig. 1g. (a) Verification on the exactly solvable $4\times 4$ case, after application of the 2-Trotter step light pulse, before any further relaxation. The exact nearest-neighbour $\Delta_{\eta}$ correlation is obtained by translating the circuit to a 32-qubit Jordan-Wigner circuit and computing the exact statevector result. Compatible with the findings in [haghshenas_digital_2025], (extrapolated) results from circuit-DMRG are reliable when a simulation fidelity $F$ of order 1 can be achieved, and are essentially random for $F\ll 50\%$. (b) Simulation fidelities on the $6\times 6$ circuits on 90 qubits, right after the light pulse (2 steps) and after up to two Hubbard-Trotter relaxation steps (3 and 4 steps), compared with the “intermediate temperature” ($\Delta\theta=2\pi/9$) quench circuit with $s=20$ Ising-Trotter steps in [haghshenas_digital_2025]. (c) Circuit fidelities on 72-qubit circuits that are obtained by deleting all gates on the ancilla qubits, as well as any maximum angle CZZ gates, which are responsible for implementing the fermion-to-qubit encoding. The result is a lower bound on the cost of natively fermionic circuit-DMRG.
@@
We now study the simulability of the circuits with Sparse Pauli Decomposition (SPD). This is a promising simulation technique that has been attracting attention recently [beguvsic2025real, broers2025scalable, lin2025utility, gharibyan2025practical, rudolph2025pauli] and which we have found most practical for the development and testing of quantum circuits in this work. Indeed, it displays a fast and precise convergence for observables with few enough gates in their light-cone for an arbitrary number of qubits (and in general, for circuits with finite magic [dowling2025magic, dowling2025bridging]). This simulation technique works in the Heisenberg picture, and consists in expanding the observable to be measured into strings of Pauli matrices
@@
A generic operator to be decomposed as in (S111) would require $4^{N}$ terms, which becomes intractable for $N\gtrapprox 20$. Starting from a local observable that involves only a few Pauli matrices, we can keep perform the simulation by storing only the non-zero terms in the decomposition (S113). However, the number of gates to keep track of grows exponentially (generically, is multiplied by $2$) after every gate application. The exact bookkeeping of all the non-zero terms ends up thus being unfeasible. As a result, different approximations can be made to get a certain precision on the result. The approximation that we will consider consist in systematically truncating, after every application of gates, all the terms $c_{n}$ that are below some threshold $\epsilon$ in absolute value. Namely, one retains in the sum (S113) only the terms $c_{n}$ such that $|c_{n}|\geq\epsilon$. The complexity of the simulation is related to the non-stabilizerness of the operator, which can be quantified by the operator stabilizer entropy [dowling2025magic]. This simulation technique is particularly appealing for finite-magic circuits and for code testing, as it is not limited by the number of qubits and the circuit geometry. This truncation scheme has been used in the largest-scale SPD numerical simulation to date [broers2025scalable].
@@
In this resource estimation study, we will first focus on the memory requirement, namely on the number of Pauli strings required in the SPD to obtain a certain precision on the observable. We will then discuss an estimation of the actual runtime based on comparison with previous works. We note that Majorana fermion decomposition (see Ref. [miller2025simulation]) instead of Pauli string decompositions are not expected to give an advantage in terms of memory, because within the fermionic encoding there should be a one-to-one correspondence between Pauli strings and Majorana strings.
@@
This is a rather standard metric for estimating the difficulty of the simulation, that has been employed before [gharibyan2025practical]. However, in our case, the circuit has a particular structure. Before the “injection”, the gates at the beginning of the circuit only act on the first $L=36$ qubits, for $90$ qubits in total. This means that in the Heisenberg picture, after the injection is passed, one can keep in the SPD only the strings that have either $I$ or $Z$ Pauli matrices on all the remaining $54$ qubits. This significantly reduces the number of strings to keep track of, but this also modifies the norm, because this projection is non-unitary. Since we would like to know the norm at the end of the circuit, we do not perform this projection. Therefore, for a fairer resource estimate, we will consider the maximal number of strings only before (in the Heisenberg picture) the injection, namely
@@
Figure S25: Relative norm squared as a function of maximal number of strings before injection $\mathcal{S}_{\rm metric}$. Data shown is for the circuits run on hardware on a $6\times 6$ lattice, performing 22 Trotter steps after state preparation (corresponding to the light pulse), 33 Trotter steps (one extra step on top of the light pulse), and 44 Trotter steps (two extra steps). The dashed lines indicate quadratic extrapolations using the last 88 points, and dotted lines linear extrapolations using the last 33 points. The gray shaded area indicates the region that is beyond the largest SPD simulation that has been performed before the present work on a supercomputer [broers2025scalable]. The blue shaded area indicates the region where $90\%$ of the norm is obtained.
@@
How do these numbers translate to actual runtimes? This question is generally difficult to answer due to differences in classical compute hardware and implementation, although in this case we can make some estimates based on the recent parallelised implementation of SPD on the Fugaku supercomputer [broers2025scalable]. In Fig. 3 of that work, wall time per gate is shown for different numbers of Pauli strings $\mathcal{S}$ and parallel processes $M$. Using plot digitisation software [WebPlotDigitizer], we have extracted the scaling of wall time per gate $w$ as a function of $\mathcal{S}$ for a fixed number of parallel processes $M=2^{7}$. While we found the slope $w\sim\mathcal{S}^{1.05}$ to fit the data best, let us make the optimistic assumption that $w\propto\mathcal{S}$ scales linear, that memory is unlimited, and that furthermore the problem can be parallelised perfectly up to $M=2^{22}$ parallel processes. Given that in that work up to 12,288 nodes were used to carry out $2^{18}$ parallel processes, this larger $M$ upper bounds the number of parallel processes that could be run on all 158,976 of Fugaku’s nodes. Under these assumptions, we find a best fit
@@
Figure S26: Estimating the run time for Sparse Pauli Decomposition on Fugaku. (a): Wall time per gate extracted from Fig. 3 in Broers ’25 which refers to Ref. [broers2025scalable] as a function of the number of Pauli strings $\mathcal{S}$. Assuming a perfectly linear scaling of wall time per gate in $\mathcal{S}$ and perfect parallelisation, i.e. inverse linear scaling in the number of parallel processes $M$, the best fit we obtain is $\mathcal{S}/M\times 100$ nanoseconds per gate. (b): Wall time per gate using $M=2^{22}$ parallel processes assuming the scaling extracted in (a) for simulating the 3-step circuit in three different scenarios: A quadratic and linear scenario, as well as the geometric mean, denoted “mid”. We consider the mid scenario to be the most likely of the three and $M^{22}$ to upper bound the maximum number of parallel processes achievable on Fugaku based on Appendix F of [broers2025scalable].
@@
Projected-entangled-pair state (PEPS) methods have become a cornerstone tensor-network framework for simulating quantum many-body systems [cirac2021matrix]. In contrast to MPS, PEPS employs a lattice-based architecture that more naturally and efficiently represents entanglement in higher-dimensional ($D>1$) systems. This enhanced expressiveness, however, comes at the cost of greater computational complexity and the absence of a canonical form for optimal compression. We employ a PEPS specifically adapted to the Octagon lattice geometry with periodic boundary conditions (PBC) and integrate a Belief Propagation (BP) algorithm to efficiently evolve the state [AlAr21].
@@
For comparison, we also report the fidelities [see Eq. (S110)] using MPS simulations, as well as those obtained using matrix-product operator (MPO) Heisenberg evolution. Notably, MPO evolution proceeds backward in time, and its accuracy depends on both the spatial support of the initial (unevolved) operator and the light-cone structure of the circuit. The fidelity of MPO simulations drops rapidly as the light cone of the operator expands, indicating that MPO methods alone are clearly incapable of simulating the full circuit. Similar to both the MPS and SPD approach, linear extrapolation of $\langle O(F)\rangle$ using MPOs has also been considered to extend the limit of finite bond dimension. Empirically, fidelity-based extrapolation of expectation values has been reported to yield results consistent with exact values in the IBM quantum utility experiment [anand2023classical]. However, the expectation value as a function of fidelity is not monotonic. Even as a heuristic method, the fidelity has to be larger than a certain threshold $\mathcal{F}_{\textrm{th}}$ to obtain meaningful signal, either by extrapolation or claiming convergence of the method. Since the MPO and SPD are similar techniques—both track the evolution of observables in the Heisenberg picture and differ only in their truncation schemes—we expect this threshold to be close to $1$, which makes the stand-alone MPO approach practically infeasible and motivates the use of hybrid techniques.
@@
One possible direction to overcome these limitations is a hybrid approach that partially evolves the state in the Schrödinger picture and the operator in the Heisenberg picture, meeting in the middle — a strategy referred to as the mixed representation. For the $4\times 4$ system, one can imagine evolving MPS and MPO and stopping near where their fidelity curves cross to achieve better fidelity. However, given the near-vanishing MPS fidelity for the $6\times 6$ system, one would instead need to evolve PEPS and MPO (or potentially PEPO) and meet at their crossing point to obtain reasonable fidelities. However, measuring observables (in the form of an MPO or PEPO) for PEPS with $\chi>8$ likely requires either a (potentially sign-problematic) sampling technique or a loop-series BP expansion [EvEtAl24]. It is unclear to us whether the observable can be accurately estimated in such a mixed representation with feasible computing resources, but it would be an interesting direction for further study. Finally, we note that all of these methods could also be implemented with an intrinsically fermionic PEPS and/or PEPO representation that avoids additional overhead of simulating the fermion-to-qubit encoding, generally resulting in lower bond dimension for the relevant objects. Reliably establishing the difficulty of each of these potential classical techniques for simulating these circuits requires significant further study and is left for future work.
@@
The second reason for choosing the mixed boundary conditions is that, ultimately, we are interested in the thermodynamic limit $N\to\infty$ where the boundary conditions have no effects anymore. In that respect, periodic, antiperiodic or mixed boundary conditions are only ways of dealing with the finite system, and are physically equally relevant. One can thus choose the boundary condition that will converge the fastest to the thermodynamic limit. It is known that averaging over “twisted” boundary conditions where a phase $e^{i\varphi}$ is given to fermions that wind around the lattice, with $\varphi$ the value to average over, make expectation values of observables converge faster to the thermodynamic limit than periodic boundary conditions [lin2001twist]. In that point of view, the mixed boundary conditions can be seen as a rough approximation of this phase averaging, and is thus expected to give results that are closer to the thermodynamic limit than periodic boundary conditions.
@@
In Fig. S28a, we show how different boundary conditions approach the thermodynamic limit. While open boundary conditions (OBC), periodic boundary conditions (PBC) and antiperiodic boundary conditions (in which every hopping term across the boundary comes with an extra minus sign, APBC) systematically overestimate the ground state energy, boundary conditions with PBC in one direction and APBC in the other underestimate it. Mixed periodic boundary conditions (mPBC), which we define by averaging over the 4 settings (PBC-PBC, APBC-APBC, APBC-PBC and PBC-APBC), thus yields the smallest finite-size effects. The resulting relative error with respect to the thermodynamic limit is shown in Fig. S28b. For the $L=6$ lattice considered in this work, mixed periodic boundary conditions already show more than 8 times smaller finite-size effects. To match the finite size effects of a $6\times 6$ mixed periodic lattice, an open boundary conditions setup would require more than $40^{2}=1600$ sites. If such an OBC system was realised using the same fermion-to-qubit mapping as described here, this would increase the qubit count from 90 to $1600\times 2\times 5/4=4000$ qubits. This discrepancy increases with increasing $L$ due to the different slopes in Fig. S28b: While in principle there exist exact expressions for the finite size relative error $r_{L}=|e_{L}-e_{\infty}|/e_{\infty}$, we can obtain a simple fit using the data points from $L=32$ to $L=40$, to obtain